3. Find all eigenvalues and eigenvectors of the matrix -2 0 2 -1 0 2 0-2 1
Answer #1

A Here Given that A= - 2 -3 Now eigen values of A is det (A-1I3)=0 -2-dok 1 > *+-17 +0+2 -- 26 7? (2+)=0 2=0,0,-4 eigene valu